OpenStudy (anonymous):

a chord of a cicle is 24 in . find the radius of the circle if chord's distance from the center is : * 5 in

OpenStudy (anonymous):



OpenStudy (anonymous):

pls..answer this..

OpenStudy (anonymous):

5^2 + 12^2 = r^2 r = 13

OpenStudy (anonymous):

what about 9 in?..

OpenStudy (anonymous):

15

OpenStudy (anonymous):

14 in?..

OpenStudy (anonymous):

4 in?...

OpenStudy (anonymous):

4^2 + 12^2 = r^2 solve for r

OpenStudy (anonymous):

10?...

OpenStudy (anonymous):

same idea

OpenStudy (anonymous):

how about \[2\sqrt{3} \]in

OpenStudy (anonymous):

12^2 + (2sqrt(3))^2 = r^2

OpenStudy (anonymous):

tnx..